OpenStudy (anonymous):

WILL FAN AND MEDAL What is the equation in point-slope form of the line passing through (-3, -4) and (0, 2)

imqwerty (imqwerty):

if you have 2 points say \((x_1,y_1)\) and \((x_2,y_2)\) then the equation of line passing through them is given by- \(\huge \frac{y-y_1}{y_2-y_1}=\frac{x-x_1}{x_2-x_1}\) now you have the 2 points just substitute the values to get the answer :)

OpenStudy (anonymous):

can you tell me which answer of these is correct and guide me through it? a. (y - 8) = 2(x - 3) b. (y + 3) = -2(x + 4) c. (y - 2) = -2(x + 4) d. (y + 4) = 2(x + 3) i'm horrible at replacing variables :(

OpenStudy (anonymous):

@imqwerty sorry for the bother hah

imqwerty (imqwerty):

okay so we have our 2 points-> \(\Large (\underbrace{-3},~\underbrace{-4})\) and \(\Large(\underbrace{0},~\underbrace{-3})\) \(x_1\) \(y_1\) \(x_2\) \(y_2\) now we can easily see \(x_1,~x_2,~y_1~and~y_2\) now we just have to substitute these values in that equation to get the answer